OpenStudy (anonymous):

if a triangle has side lengths of 7,4,and 10, what kind of triangle is it? (acute,right,or obtuse)

OpenStudy (anonymous):

To test for right triangle: \[a^2 + b^2 = c^2 \] \[4^2 + 7^2 = 10^2 \] 16 + 49 = 100 65 = 100 FALSE Therefore, it is not a right triangle.

OpenStudy (anonymous):

so than it is an acute than?

OpenStudy (anonymous):

\(a^2 + b^2 = c^2\) -- right triangle \(a^2 + b^2 > c^2\) -- acute-angled \(a^2 + b^2 < c^2\) -- obtuse-angled

OpenStudy (anonymous):

oh its obtuse

OpenStudy (anonymous):

Precisely :)

OpenStudy (anonymous):

oh okay thank you! :)

OpenStudy (anonymous):

yw :)